OpenStudy (anonymous):

Locate all inflection points of:

OpenStudy (anonymous):

well lets see this dont make sence so i cant help sorry i wish i could

OpenStudy (anonymous):

What's the function you're looking at?

OpenStudy (anonymous):

\[ f(x) = x ^{4} + 6x ^{3} - 24x ^{2} + 26\] I get the answer of x=1 and x=-4

OpenStudy (anonymous):

This is a calculus class?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh good. Didn't want to try to factor that ;p

OpenStudy (anonymous):

i did not know what she wanted because im in the 6th grade haha

OpenStudy (anonymous):

\[f'(x) = 4x^3 + 18x^2-48x = 0\] I'm getting \(x \in \{1.8807, 0, -6.3807\} \) for the critical points.

OpenStudy (anonymous):

this looks so confusing hehe

OpenStudy (anonymous):

And of those, the only inflection point is the 0.

OpenStudy (anonymous):

but i think ill get it soon when i get in a different grade

OpenStudy (anonymous):

It's ok moon, keep practicing and you'll get to calculus sooner than you think.

OpenStudy (anonymous):

did you take the 2nd derivative?

OpenStudy (anonymous):

ok thanks polpak

OpenStudy (anonymous):

i became a fan of you

OpenStudy (anonymous):

\[f''(x) = 12x^2+36x - 48 = 0\] \[\implies x^2 + 3x - 4 = 0 \implies x \in \{1,-4\}\] So actually 0 isn't an inflection point either.

OpenStudy (anonymous):

Since none of the critical points are also points at which f'' is 0 you don't have any points of inflection.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

from what i am reading in my book the inflection points is a point where the concavity changes. when i test some points i come up with\[x <-4 = positive \]\ \[x > 1 = pos\]\[-4 < x < 1 = negative \] so it looks like it changes from positive at -4 to negative then postive again at 1

OpenStudy (anonymous):

That is what I get, too.

OpenStudy (anonymous):

Yes that's true. Technically they are inflection points. I just typically think of inflection points as in the stationary kind. Those are non-stationary points of inflection because while the concavity does change, the tangent to the function is non-zero.

OpenStudy (anonymous):

ok for a question on is problem I need to"show inflections point(s) and explain in complete sentence why these number are inflection points and not critical points".

OpenStudy (anonymous):

So I am kinds confussed with this

OpenStudy (anonymous):

"kinda"

OpenStudy (anonymous):

They are inflection points because the curvature changes (goes from positive to negative and then negative to positive). They are not critical points because the slope of the tangent to the curve at those points is not 0.

OpenStudy (anonymous):

i see now. thanks!!